OpenStudy (anonymous):

Write an equation of a line that passes through the point (-5, 7) and whose slope is 6. Answer y + 5 =6(x + 7) y + 7 = 6(x -5) y + 7 = 6(x +5) y – 7 = 6(x +5)

OpenStudy (anonymous):

@jayz657 @privetek

OpenStudy (anonymous):

@emilie.li

OpenStudy (anonymous):

Help

OpenStudy (anonymous):

@EpikRebel HELPP

OpenStudy (anonymous):

y-y1=m(x-x1) y-7=6(x-(-5)) y-7=6(x+5)